Two 2.3-cm-diameter-disks spaced 1.9 mm apart form a parallel-plate capacitor. The electric field between the disks is 5.0×105 V/m . How much charge is on each disk? b. An electron is launched from the negative plate. It strikes the positive plate at a speed of 2.2×107 m/s . What was the electron's speed as it left the negative plate?
Answer #1

Voltage, V = Ed = (5 x 10^5) x 0.0019 = 950 V

Charge, q = (Ae0/d) V = (pi x 0.0115^2 x 8.85 x 10^-12/0.0019) x 950 = 1.84 x 10^-9 C

B)

Force, F = qE = (1.6 x 10^-19)(5 x 10^5) = 8 x 10^-14 N

Acceleration a = F/m = (8 x 10^-14)/(9.11 x 10^-31) = 8.78 x 10^16 m/s^2

V^2 = u^2 + 2as

V^2 = (2.2 x 10^7)^2 - 2 x (8.78 x 10^16) x 0.0019

V = 1.23 x 10^7 m/s

Comment in case any doubt please rate my answer ...